OpenStudy (anonymous):

Systems of Equations in three variables Determine whether (-3, -1, 4) is a solution of the system 2x-2y-3z=-16 4x-3y+3z= 3 x+ y - 5z = -24 I'm having troubles with this problem could someone explain this problem to me. Also this next problem as well please... solve the system. x+ y+ z = 7 x- y= 2z = 4 2x+ y + z = 5

OpenStudy (anonymous):

thank you @JamesWolf

OpenStudy (anonymous):

well, as far as determining if a point is a solution. Whack it into every equation and see if they hold to solve the system, you need to get an equation for just one of the variables. Theres loads of ways you could do this, what i would do is add equation 1 and 2 to get x + y + z + x - y + 2z = 7 + 4 2x + 2z = 11 then I would take away 2x the top row from this 2x + 2z - 2x -2y -2z = 11 -7 2y = 4 y = 2. now you have a way in. you can replace this value of y into any of the equations and solve for x or z. replace that into "2x + 2z = 11" to solve for an other variable. and so on like this until you get all the variables. generally hteres no quick way to do it, and its a matter of doing a few to spot which moves will solve it quickest

OpenStudy (anonymous):

im presuming the second line was not x- y= 2z = 4 but x - y + 2z = 4 ?

OpenStudy (anonymous):

well thats wrong somewhere lol. http://math.bd.psu.edu/~jpp4/finitemath/3x3solver.html using that to check the answer

OpenStudy (anonymous):

yes the = was suppose to be a +

OpenStudy (anonymous):

that solves them and gives you steps

OpenStudy (anonymous):

click the little + by the show steps, to see either elimination method or substitution

OpenStudy (anonymous):

Thank you!!! This helps a lot thank you yea the fact that it will show its steps helps because I am a very visual person.

OpenStudy (anonymous):

Yeah me too. Nice that it gives you both elimination or substitution method

OpenStudy (anonymous):

Yes definitely do you have any more math sites? That you have that help you?